OpenStudy (anonymous):

M is the midpoint of JK. The coordinates of J are (6, 3) and the coordinates of M are (-3, 4), find the coordinates of K.

OpenStudy (anonymous):

Use the midpoint formula" http://geometry.jdeer.com/Images/c1/midpoint%20formula.gif

OpenStudy (***[isuru]***):

the coordinates of the midpoint (c) can be find by this formula \[c = [ \frac{ mx_{1} + nx_{2} }{ m +n } , \frac{ my_{1} + ny_{2} }{ m + n } ]\] where where x1 , y1 will be the coordinates of the point J and x2 and y2 will be the coordinates of point K and m and n is the ratio between the distance between JM and MK got it so far ?

OpenStudy (anonymous):

yea im understanding what your doing :)

OpenStudy (***[isuru]***):

since the M is the midpoint of JK , the ratio of JM to MK will be 1 ( because JM =MK) then n= 1 and m = 1 then we can simplify the equation in to \[c = [\frac{ x_{1} + x_{2} }{ 2 } , \frac{ y_{1} +y_{2} }{ 2 } ]\] ok ?

OpenStudy (***[isuru]***):

and we already know the coordinates of c ,therefor... \[( -3 , 4 ) = [ \frac{ x_{1} + x_{2} }{ 2} , \frac{ y_{1} + y_{2} }{ 2 } ]\] we also know the coordinates of J which is ( 6 , 3 ) represent by x1 and y1 , therefor \[( - 3 , 4) = [\frac{ 6 + x_{2} }{ 2 } , \frac{ 3 + y_{2} }{ 2 } ]\] now we can get 2 seperate equations from them

OpenStudy (anonymous):

still following

OpenStudy (***[isuru]***):

considering the x coordinates -3 = ( 6 + x2) /2 -6 = 6 + x2 ----(1) considering the y coordinates 4 = ( 3 + y2) /2 8 = 3 + Y2 ----(2) using (1) and (2) u can find the coordinates of the point K hope this will help ya!!

OpenStudy (***[isuru]***):

did u got it or did i make it a mess ? :)

OpenStudy (anonymous):

yea K =5,6 :)

OpenStudy (***[isuru]***):

is K = 5 , 6 is the given answer ?

OpenStudy (anonymous):

thats what i got from it (5,6)

OpenStudy (***[isuru]***):

"thats what i got from it (5,6)" "from it" u mean wt ? the equation i wrote or something else ?

OpenStudy (anonymous):

thats what i got as the answer from the equation you set up for me

OpenStudy (***[isuru]***):

mmm.... i think u need look twice... u got the x coordinate as 5 by solving -6 = 6 + x2 and y coordinate as 6 by solving 8 = 3 + Y2 try to solve them again.......

OpenStudy (anonymous):

now im confised lol

OpenStudy (***[isuru]***):

ok.... lets c where did u got it wrong... can u pls tell me the steps u follow to get x coordinates as 5 ?

OpenStudy (anonymous):

its ok, dont worry about it. thanks for trying, maybe its just me being tired

OpenStudy (***[isuru]***):

u r welcome !! :)